OpenStudy (anonymous):

Solve: -28 = -7x x = -4 x = 4 x = -196 x = 196

OpenStudy (anonymous):

4

OpenStudy (anonymous):

you want to isolate you "x" on one side of the "=" sign, and your numbers on the other first thing you want to do, is let x be by itself, so you divide each side by 7, - 28/7 = x -4 = x

OpenStudy (anonymous):

algelvr...u have to divide by -7....and it's positive 4